# Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning

### v2

@@ 4.1 Motivation
Research in Cognitive Science highlights the importance of reflection, organization, and linking information as critical components of effective learning. Studies show that when learners take time to reflect on their experiences, they deepen their understanding and enhance retention [24]. This reflective practice encourages individuals to connect new information with existing knowledge, fostering a more integrated learning experience [2]. Moreover, organizing information into coherent structures, such as outlines or concept maps, allows learners to see relationships between concepts, making it easier to retrieve information later [19]. Linking new material with relevant prior knowledge—often referred to as associative learning—further strengthens memory retention [26] and also benefits future learning.
Our Approach.
It is evident that besides reflection, comparison of old and new information, the key factor in learning efficiently is to organize and exploit them in an insightful way, where concepts are linked and arranged according to their semantic meanings. This observation motivates us to develop a deep learning classifier that learns labels or concepts structured in a hierarchical taxonomy. The aim is to enable the classifier to grasp relevant concepts more effectively, helping to mitigate the challenge of catastrophic forgetting. More specifically, we propose structuring the data labels in a hierarchical taxonomy, which can be dynamically constructed using domain expertise, adapting as needed based on the specific context and evolving understanding of the domain. Based on this structure, we establish a reference framework for the relationships between classes, identifying which classes belong to the same group with many shared characteristics, easily confused due to overlap, and require more focus (see Figures 3, 4). This approach not only helps the model better avoid forgetting, but also reinforces knowledge to facilitate future learning.
@@
Figure 4: t-SNE visualization of classes within leaf groups of Four-legged animals ($\bullet$ circular points) and Plants ($\blacktriangle$ triangular points) when learning Task 3, Split-CIFAR100.
@@
Taking the learning process of Split-CIFAR100 as an example, when training on task $t=3$, we can construct a tree-like taxonomy of concepts/labels, as shown in Figure 3. We observe that the concepts/classes under the same leaf in the tree-like taxonomy (e.g., oak tree, willow tree, and pine tree in the "plants" leaf group) exhibit stronger visual and semantic correlations. Consequently, as shown in Figure 4a, their features in the feature space become more overlapping compared to those from other leaf groups (e.g., otter and hamster under the "four-legged" leaf group), leading to more confusion and performance degradation when predicting these concepts/classes. This highlights the impact of organizing concepts/labels in a tree-like taxonomy, where data examples within the same leaf group share stronger visual and semantic relationships, causing greater overlap and increased confusion during predictions. Linking to our analysis in Section 3.2, the tree-like taxonomy of concepts/labels serves as a tool to help identify easily confused classes/concepts, facilitating the subsequent process of making them more distinct and separable in the feature space.
Our approach aims to train the backbone network so that all class representations must be distinct, especially those within each leaf group, to achieve maximum distinguishability. For the leaf group of four-legged animals, we assume that "mouse" and "porcupine" arrive in Task 1, while "otter" and "hamster" are in Task 3. When learning Task 1, to classify "mouse" and "porcupine", the backbone is encouraged to capture the essential features of four-legged mammals to efficiently differentiate between these two animals. We hope that the knowledge learned from "mouse" and "otter" in Task 1 can be beneficial for the next tasks. Then in Task 3, we again learn to distinguish "hamster", "otter" and these old ones in this group. In this way, the mechanism helps further strengthen the learning of more efficient and robust features for the Mammals group.
To summarize, by grouping and categorizing, we expect the model to concentrate more on the detailed features of each leaf group. This enhances its ability to distinguish related objects and reinforces the model’s knowledge of each group. Therefore, this strategy not only alleviates the forgetting of old knowledge—often caused by new classes that are difficult to distinguish from old ones within the same leaf group, but also enables active knowledge transfer between tasks.
4.2 Training phase
4.2.1 Exploiting hierarchical label/concept taxonomy
Figure 5: We focus on separate easily confused classes within each leaf group.
@@ 4.2.1 Exploiting hierarchical label/concept taxonomy
During the training process, whenever a new class appears, its label name is automatically added to the tree-like taxonomy, into a leaf group containing classes with similar characteristics (Figure 3). To develop this hierarchical structure, we can rely on expert knowledge, for example, ChatGPT, which can help us incrementally construct a meaningful and semantically-related tree (see Appendix B). Structuring information in this way not only aligns with how the human brain effectively connects and remembers information but also provides useful insights during training, indicating how each knowledge is related to the other and which requires further focus.
@@
As analyzed above, reflecting on and organizing knowledge is the key factor for efficient learning. That is, the model should always be encouraged to identify the decision boundary between all old and new classes, especially those in the same leaf group. Assume that we finished the task $t-1$ and are learning the prompt $\bm{P}_{t}$ for task $t$, our aim is to learn the backbone network $f_{\Phi,\bm{P}_{t}}$ that can minimize overlap between all classes so far, especially focus on increasing the separability between classes belonging to the same leaf group extracted from the taxonomy (e.g., four-legged mammals, plants, etc.,). Let $g\in\mathcal{G}$ be a leaf group, ${X}^{g}_{k}$ and ${Y}^{g}_{k}$ denote the corresponding sets of input samples and labels under the group $g$ that belong to the task $k$ ($k\leq t$). Besides Cross Entropy loss $\mathcal{L}_{CE}$, we propose using a regularization loss function for sample ${\bm{x}}$ (arrives in task $t$, belong leaf group $g$) as follows:
$\mathcal{L}_{\mathcal{G}}(\psi,\bm{P}_{t},\bm{x})=-\alpha\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}-\beta\mathcal{L}_{all},$ (1)
@@
where $\mathcal{L}_{all}=\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$ is the Supervised Contrastive loss that we force all class representations so far to separate from each other, $u({\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}})=\exp(\frac{{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}}}{\tau})$, with ${\bm{z}}_{\bm{x}}=f_{\Phi,\bm{P}_{t}}({\bm{x}})$ is the feature vector on the latent space of the prompt-based model, $y_{\bm{x}}$ is the ground truth label of ${\bm{x}}$, $\tau$ is a temperature ($\tau=0.1$ for all experimental setting), and $\alpha$ is the coefficient that controls how much we want to force on classes belonging to the same leaf group stay apart further. For each data sample ${\bm{x}}|y_{\bm{x}}\in Y_{k<t}$, the corresponding representation ${\bm{z}}_{\bm{x}}$ is sampled from the Gaussian Mixture model $\text{GMM}_{y_{\bm{x}}}=\{\mathcal{N}(\mu_{y_{\bm{x}}},\Sigma_{y_{\bm{x}}})\}_{k=1}^{K}$ of the respective class, which is obtained at the end of each corresponding task. This technique of using pseudo features of old data is also employed in many existing prompt-based CL methods, as the shift of features is minimal (Table 2).
Equation 1 implies that when learning a new task, new classes/new knowledge will be compared and contrasted with existing ones. That is, the model will be encouraged to identify the decision boundary between old and new classes, especially focusing on those in the same leaf group via controllable coefficient $\alpha$ (Figure 5). Furthermore, focusing on the specific knowledge within each leaf group helps our model strengthen and consolidate its understanding of this domain, especially when the current prompt is initialized by the previous ones. This is achieved by employing a prompt ensemble strategy similar to that in [30]: $\bm{P}_{t}=\eta\bm{P^{\prime}}_{t}+(1-\eta)\sum_{i=1}^{t-1}\bm{P^{\prime}}_{i}$, where $\left\{\bm{P^{\prime}}_{i}\right\}_{i=1}^{t}$ is the set of learnable prompt elements which are used to adapt to corresponding tasks, the prompt elements of previous tasks are kept fixed ($\eta=0.99$ for all setting).
4.2.2 An optimal transport-based approach to further exploiting priori from pretrained model
Considering a leaf group, there may be data classes with varying levels of overlap in the latent space(Figure 4, 6). Although focusing on classes within the same leaf group helps improve the ability to recognize difficult-to-identify classes, we still treat all classes in that group equally. Thus, the algorithm may inadvertently ignore important pairs of classes that are easily confused and need to be further distinguished. Besides, pretrained models are known to have been extensively trained on large datasets, resulting in a substantial repository of generalization ability. Therefore, the prior knowledge from these models often provides a valuable starting point for the adaptation to downstream tasks. However, we seem to frequently overlook the initial behavior of pretrained models on the training data, particularly regarding relations between classes, which classes are easily classified and which are prone to confusion.
@@ 4.2.2 An optimal transport-based approach to further exploiting priori from pretrained model
Figure 6: Wasserterin distance between classes (Split-CIFAR100) in latent space of pre-trained backbone (Sup-21K).
Therefore, in this work, we propose to exploit the pre-trained model from a new perspective, which can comprehend the use of the label-based tree-like taxonomy during training, where we can take advantage of prior assumptions about the relationships between the image classes. Firstly, to extract the relationship between the classes, we use L2 Wasserstein distance (WD) to compare the distributions of feature vectors of each pair of class. In particular, let $D_{c_{i}}^{\Phi}$ be the distribution of class $c_{i}$ on the latent space of the pretrained model $f_{\Phi}$, which is obtained in the form of a Gaussian Mixture model at the end of the respective task. When $t$ tasks have arrived, we have the corresponding sets of distributions $\{D_{c}^{\Phi}\}_{c\in Y_{1,t}}$ of all $m_{t}$ classes from all tasks so far. Therefore, we gradually complete the WD-based matrix between pairs of classes:
$M=[W_{2}(D_{c_{i}}^{\Phi},D_{c_{j}}^{\Phi})]_{m_{t}\times m_{t}}.$ (2)
We then compute the weight matrix $\Gamma=\left[\gamma_{ij}\right]_{m_{t}\times m_{t}}=\left[1/{\exp(M_{ij}/\delta)}\right]_{m_{t}\times m_{t}}$, where $\delta$ is a temperature. We then apply this information to obtain a weighted version of $\mathcal{L}_{\mathcal{G}}$, in which the closer the two class distributions are, the larger the weight assigned, and they will be focused to push away. Consequently, our regularization loss becomes:
$\mathcal{L}_{\mathcal{G}}(\psi,\bm{P}_{t},\bm{x})=-\alpha\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}\gamma_{y_{x}y_{\tilde{x}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}-\beta\mathcal{L}_{all}.$ (3)
This strategy is completely economical and aligns well with the CL learning scheme as matric $M$ is continuously expanded and provides useful information for training new tasks. Practically, when learning a new task, the first epoch is spent capturing information about the behavior of the pretrained model on the data for this task. Moreover, this approach is similar to the findings in Cognitive Science [22, 1], showing that the accumulated experiences from past learning create momentum for learning new skills more effectively.
4.3 Testing phase
We observe that classes within the same leaf group often share many common characteristics. Therefore, we propose a testing strategy that leverages information from each leaf group to gain a new perspective on the identification of data samples, especially those at the boundaries between different leaf groups.
In particular, the final prediction for sample $\bm{x}$ will be modified based on the probability that $\bm{x}$ belongs to a certain leaf group $g\in\mathcal{G}$ (i.e., $p(g|\bm{x},\bm{P}_{\hat{t}})$). Intuitively, if the representation $\bm{z}_{x}$ of $\bm{x}$ has many similarities with class representations in group $g$, then $p(g|\bm{x},\bm{P}_{\hat{t}})$ will increase, thereby raising the likelihood that $\bm{x}$ belongs to the corresponding classes in $g$ (i.e., $p(y|g,\bm{x},\bm{P}_{\hat{t}})$):
$p(y|\bm{x})=p(y|\bm{x},\bm{P}_{\hat{t}})=\sum_{g\in\mathcal{G}}p(g|\bm{x},\bm{P}_{\hat{t}})\cdot p(y|g,\bm{x},\bm{P}_{\hat{t}})=\sum_{g\in\mathcal{G}}p(g|\bm{x},\bm{P}_{\hat{t}})\cdot p(y|x,\bm{P}_{\hat{t}})\cdot\mathbb{I}_{y\in Y^{g}},$ (4)
where $\mathbb{I}_{y\in Y^{g}}=1$ if $y\in Y^{g}$, else 0. The value of $p(g|\bm{x},\bm{P}_{\hat{t}})$ is calculated based on the "energy" of $\bm{x}$ w.r.t group $g$, in relation to other group $g^{\prime}\neq g$:
$p(g|\bm{x},\bm{P}_{\hat{t}})=\frac{\exp\{E(\bm{x},g)\}}{\sum_{g^{\prime}\in\mathcal{G}}\exp\{E\left(\bm{x},g^{\prime}\right)\}}.$ (5)
In Eq. (5), $E\left(\bm{x},g\right)$ indicates the "energy" of $\bm{x}$ w.r.t leaf group $g\in\mathcal{G}$. Remind that for each class $c$, we maintain a GMM of $K$ mixtures $\{\mathcal{N}(\bm{\mu}_{c,i},\Sigma_{c,i})\}_{i=1}^{K}$. Based on the prototypes for a class, we can define the distance from $\bm{x}$ to a class $c$ as $d({\bm{x}},c)=\min_{1\leq i\leq K}cosine\_distance({\bm{z}}_{x},\bm{\mu}_{c,i})$. Limiting to the group $g$, we define $\hat{y}_{\bm{x}}^{g}=\operatorname*{arg\,min}_{c\in Y^{g}}d({\bm{x}},c)$ (i.e., $Y^{g}$ is the set of all classes in $g$). We define the energy of interest as
$E(\bm{x},g)=-d({\bm{x}},\hat{y}_{x}^{g})-\xi\sum_{c\in Y^{g}}\gamma_{c,\hat{y}_{x}^{g}}\sum_{i=1}^{K}\sqrt{(\bm{z_{x}}-\mu_{c,i})^{T}\Sigma^{-1}_{c,i}(\bm{z_{x}}-\mu_{c,i})}$ (6)
where the first terms is the cosine similarity between $\bm{z_{x}}$ and the closest class prototype within group $g$ and $\gamma_{c,\hat{y}_{x}^{g}}$ is the value obtained from the weight matrix $\Gamma$ (Section 4.2.2) - indicating the correlation between class $c$. This approach exploits the correlation between ${\bm{x}}$ and $g$ while reducing the disadvantage of large groups with many classes, whereby the distances of classes that are less related to $y$ will have less weight and vice versa. Finally, $\xi$ is the hyperparameter, which controls the amount of information referenced from the group.
By this strategy, features $\bm{z_{x}}$ will have an additional point of view to determine which class that $\bm{x}$ is more likely to belong to, especially for $\bm{z_{x}}$ located at the boundary between leaf groups—such as between the group of "Plants" and the group of "Four-legged animals". This is similar to how having more prior knowledge improves posterior probability in Bayes’ rule and how humans with more in-depth knowledge in different fields have greater experience in solving problems.
5 Experiments
5.1 Experimental setup
Datasets. We examine widely used CIL benchmarks, including Split CIFAR-100, Split ImageNet-R, 5-Datasets, and Split CUB-200 (please refer Appendix A.1 for more details).
Baselines. We compare our method with notable CL methods exploiting prompt-based approach for pre-trained models, including the methods using shared prompts for all tasks: L2P [35], DualPrompt [34], OVOR [9]; and the methods dedicates a distinct prompt set for each task like: S-Prompt++ [33], CODA-Prompt [25], HiDe-Prompt [30], CPP [15].
Metrics. We use two main metrics, including the Final Average Accuracy (FAA), denoting the average accuracy after learning the last task, and the Final Forgetting Measure (FFM) showing the forgetting of all tasks after learning the sequence of tasks (see Appendices A.2 & A.3).
The implementation is described in detail in Appendix A.4.
5.2 Experimental result
Our approach achieves superior results compared to baselines.
Table 2 presents the overall performance comparison between our proposed method and all the baselines. The key observation is that our method is the strongest one with the gap between our method and the runner-up method is about 2% in terms of FAA on all considered datasets. Additionally, the results show that our method avoids forgetting better than all baselines, notably reducing forgetting by more than 2% on the Split-CIFAR100 dataset compared to the strongest one.

### v3

@@ 4.1 Motivation
Research in Cognitive Science highlights the importance of reflection, and organization of information as critical components for effective learning. Studies show that when learners take time to reflect on their experiences, they deepen their understanding and enhance retention [37]. This reflective practice encourages individuals to connect new information with existing knowledge, fostering a more integrated learning experience [3]. Linking new material with relevant prior knowledge—often referred to as associative learning—further strengthens memory retention [39] and also benefits future learning. Moreover, organizing information into coherent structures, such as outlines or concept maps, allows learners to see relationships between concepts, making it easier to retrieve the information later [30].
@@
Figure 4: t-SNE visualization of classes within leaf groups of Four-legged animals ($\bullet$ circular points) and Plants ($\blacktriangle$ triangular points) when learning Task 3, Split-CIFAR100. The order of appearance of the classes: Task 1 - ”mouse”, ”porcupine”, ”oak tree”; Task 2 - ”willow tree”; Task 3 - ”otter”, ”hamster”, ”pine tree” (Figure 2 and 3). We can see that if we train tasks independently with $\mathcal{L}_{CE}$, the classes within each such group, which arrive at different tasks, can be overlapped seriously (Figure 4(a)).
Our Approach.
It is evident that besides reflection - comparison of old and new knowledge, the key factor in learning efficiently is to organize and link information in an insightful way, where concepts are linked and arranged according to their semantic meanings. This observation motivates us to develop deep learning models that learn concepts structured in a hierarchical taxonomy. The aim is to enable the models to grasp relevant concepts more effectively, thereby enhancing distinguishability and mitigating catastrophic forgetting. More specifically, we propose structuring the data labels in a tree-like taxonomy, which can be flexibly expanded over time using domain expertise. Based on this structure, we have a view of the relationships between classes, identifying which classes belong to the group with many shared characteristics, easily confused, and require more focus (see Figures 3, 4).
@@
Taking the learning process of Split-CIFAR100 (see Appendix A.1) as an example, when training on task $t=3$, we can construct a tree-like taxonomy of concepts/labels, as shown in Figure 3. We observe that the classes under a same leaf in the structure (e.g., oak tree, willow tree, and pine tree in the leaf group ”plants”) exhibit stronger visual and semantic correlations. Consequently, as shown in Figure 4a, these class features are more overlapping with each other rather than with another class from other leaf groups (e.g., ”four-legged”). In this way, the tree-like taxonomy serves as a tool to help identify easily confused classes, facilitating the subsequent process of making them more separable in the feature space. Leveraging the insight obtained from the taxonomy above, we aim to train a network so that all class representations must be distinct, especially those within each leaf group. Over time, this strategy not only helps the model better mitigate forgetting caused by representation overlap but also reinforces existing knowledge to facilitate future learning.
4.2 Leveraging Hierarchical Taxonomies for Better Representation Learning
@@
During the training process, whenever a new class appears, its label name is automatically added to the tree-like taxonomy, into a leaf group containing classes with similar characteristics (Figure 3). To develop this hierarchical structure, we can rely on expert knowledge, for example, ChatGPT, which can help us incrementally construct a meaningful and semantically related tree (see Appendix B). Structuring information in this way not only aligns with how the human brain effectively connects and remembers information but also provides useful insights during training, indicating how each knowledge is related to the other and which requires further focus.
@@
For each task $t$, we dedicate a set of prompt $\bm{P}_{t}$. In each task, we aim to minimize overlap between all classes so far, especially focusing on increasing the separability between classes belonging to the same leaf group on the taxonomy (e.g., four-legged mammals, plants, etc.,). That is, taxonomy acts as a reference information channel to support the training process. Let $g$ be a leaf group on the taxonomy $\mathcal{G}$ (i.e., $g\in\mathcal{G}$), ${X}^{g}_{k}$ and ${Y}^{g}_{k}$ denote the corresponding sets of input samples and labels under the group $g$ that belong to the task $k$ ($k\leq t$). Besides Cross Entropy loss $\mathcal{L}_{CE}$, we propose using a regularization loss function for sample ${\bm{x}}$ that arrives in task $t$ and belongs to leaf group $g$ as follows:
$\mathcal{L}_{\mathcal{G}}(\psi,\bm{P}_{t},\bm{x})=\alpha\mathcal{L}_{g}(\psi,\bm{P}_{t},\bm{x})+\beta\mathcal{L}_{all}(\psi,\bm{P}_{t},\bm{x})$ (1)
where we have defined
$\mathcal{L}_{g}(\psi,\bm{P}_{t},\bm{x})$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
$\displaystyle\mathcal{L}_{all}(\psi,\bm{P}_{t},\bm{x})$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
@@
are the Supervised Contrastive loss that we put on class representation within leaf group $g$, and all classes so far, respectively; $u({\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}})=\exp(\frac{{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}}}{\tau})$, with ${\bm{z}}_{\bm{x}}=f_{\Phi,\bm{P}_{t}}({\bm{x}})$ is the feature vector on the latent space of the prompt-based model and $y_{\bm{x}}$ is the ground truth label of ${\bm{x}}$, $\tau$ is the temperature with $\tau=0.1$ for all experiments, and $\alpha$ is the coefficient that controls how much we want to force classes within leaf group $g$ stay apart further.
In addition, because the data of old tasks can not be accessed when training the new one, we encode the information of each trained class $c$ in the form of a Gaussian Mixture model $\text{GMM}_{c}=\{\mathcal{N}(\mu_{c,i},\Sigma_{c,i})\}_{i=1}^{K}$, at the end of the task, where $K$ is the number of components. Then in task $t$, the representation of each old class $c$ is sampled from $\text{GMM}_{c}$, to compute loss functions in Eq. (1), supporting the training of the current task.
Looking closer at Eq. (1), this equation implies that when learning a new task, the model will be encouraged to identify the decision boundary between old and new classes, especially focusing on those in the same leaf group (see Figure 2). This manipulation matches the way our brain naturally works - without reflection and comparison, we would be confused about old and new concepts, especially those sharing many common characteristics, leading to incorrect judgments and decisions in practice. For A.I models, learning new tasks without thoroughly considering the learned information of old tasks can lead to uncontrolled overlapping of layer representations in the latent space, resulting in forgetting and harming final performance. In this work, thanks to the insight from the hierarchical taxonomy, we can identify which group of classes that easy to get overlapped in the latent space, thereby actively intervening and enhancing the model representation learning.
@@
Figure 5: Wasserterin distance between classes (Split-CIFAR100) in latent space of pretrained backbone (Sup-21K).
4.3 Harnessing Prior Knowledge from Pretrained Models via Optimal Transport
Considering a leaf group, there may be data classes with varying levels of overlap in the latent space. For example, Figure 5 provides statistical results regarding the relative position of classes in leaf groups ”plants” and ”four-legged mammals”. We can see that in the group of animals, the L2-Wasserstein distance between class representations of ”mouse” and ”otter” (39.30) is more significant than the one between ”mouse” and ”hamster” (26.05), meaning that the class representation of the second pair can overlap more than the first one. Although focusing on classes within the same leaf group helps improve the ability to recognize difficult-to-identify classes, we still treat all classes in that group equally. Thus, the algorithm may inadvertently ignore the pairs of classes that are easily confused and need to be further distinguished.
Moreover, pretrained models are extensively trained on large datasets, which results in their substantial generalization abilities. That is the reason why they are considered as good starting points for the adaptation to downstream tasks [8, 4, 17, 15, 35]. In this work, we propose to exploit the pretrained models from a novel perspective, which can comprehend the use of the label-based tree-like taxonomy during training. In particular, we take advantage of these models to obtain prior assumptions about the relationships between the class representations on the latent spaces, thereby elegantly intervening in better separating each pair of classes within a leaf group.